# Revisiting Multivariate Time Series Forecasting with Missing Values

### v2

@@ 3.5 Consistency Regularization
While the IB framework encourages learning a compact representation, high missing rates can still lead to unstable training as shown in Sec. E.2, where the model overfits to the specific variate in a given time window (Choi & Lee, 2023). To mitigate this and enhance robustness, we introduce a consistency regularization scheme (Bachman et al., 2014; Laine & Aila, 2016). The core intuition is that the model’s prediction should be invariant to the missingness. We achieve this by creating an augmented, more challenging view of the input, e.g, introducing additional noise to partially observed data. By enforcing that the representations learned from the observed and augmented views remain consistent, we regularize the model to handle missing values while stabilizing the refined representations instead of focusing excessively on a limited subset of observed data and neglecting crucial task-relevant variate correlations.
@@ 3.6 Model Learning
where $\alpha,\beta,\gamma\in\mathbb{R}$ are the preset balancing coefficients. This entire guidance helps CRIB extract the most important task-relevant information from the partially observed time series data while filtering out irrelevant noise introduced by missing values.
4 Experiment
In this section, extensive experiments on four real-world time series forecasting datasets are conducted to illustrate the effectiveness of our proposed CRIB. More experiments are in Appendix E.
4.1 Experiment Settings
Datasets.
We evaluate our model on four MTSF datasets: PEMS-BAY (Li et al., 2017), Metr-LA (Li et al., 2017), ETTh1 (Zhou et al., 2021), and Electricity (Wu et al., 2021). The key statistics and information of these datasets are summarized in Appendix B. To assess the model’s effectiveness and robustness in handling missing values, we introduce synthetic missingness by randomly removing data points at varying missing rates of 20%, 40%, 60%, and 70% with three different missing patterns. During the experiments, we normalized the data to facilitate better model fitting.
Baselines.
We chose 12 representative models for performance comparison. (1) Representative MTSF-M methods: BRITS (Cao et al., 2018), SAITS (Du et al., 2023), SPIN (Marisca et al., 2022), GRIN (Cini et al., 2021), and BiTGraph (Chen et al., 2023). (2) Transformer-based MTSF methods: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), and PAttn (Tan et al., 2024). (3) MLP-based and RNN-based MTSF methods: DLinear (Zeng et al., 2023), WPMixer (Murad et al., 2025), TimeXer (Wang et al., 2024b), and SegRNN (Lin et al., 2023).
Since the last two kinds of methods are not designed for MTSF-M tasks, we also study their variants by combining them with the current SOTA time series imputation method TimesNet (Wu et al., 2022) to build a two-stage framework, where TimesNet imputes and they predict. To simulate a practical scenario where the ground truth for missing values is unavailable during inference, TimesNet is trained on each dataset with a 10% missing rate and then imputes the observed data with 20%, 40%, 60%, and 70% missing rates. The original models and the variants are denoted as Original and Imputed, respectively. More baseline details are in Appendix C.
Implementation Details.
We use Adam optimizer (Kingma, 2014) to learn the parameters of all models with $10^{-3}$ learning rate. The unified-variate attention of CRIB is configured with 2 layers and 4 heads, while the predictor is implemented as a simple 2-layer MLP. Both historical and future time window sizes are set to 24 for all methods, following the setting of BiTGraph (Chen et al., 2023). The patch length is set to 8, so every time series in a time window is patched into three tokens. The entire dataset is divided into training, validation, and testing sets with ratios of 60%, 20%, and 20%. Hyperparameters of all baselines are consistent with their original papers.
Metrics. In our experiments, we use Mean Absolute Error (MAE) and Mean Squared Error (MSE) to evaluate the forecasting performance of different methods.
4.2 Main Results
@@
The average performance comparisons between baselines and CRIB across four datasets are presented in Tab. 1, with full results in Appendix D and more missing patterns performance comparison in Figs. 3 and E.3. We denote out-of-memory and improvement as OOM and IMP, respectively. Based on these results, we summarize our observations (Obs.) as follows:
Obs. ❶: CRIB demonstrates superior performance improvement in MTSF-M tasks. As shown in Tabs. 1, 4, 3 and E.3, CRIB achieves the lowest MAE and MSE across all 4 datasets and 3 missing patterns, with substantial improvements. Specifically, CRIB reduces the MAE by over 18% on ETTh1 and over 13% on PEMS-BAY compared to the strongest baseline. We attribute this improvement to our model’s design, which integrates patch embedding, unified-variate attention, and consistency regularization under the IB principle, thus enabling CRIB to effectively filter noise from incomplete data while preserving essential predictive signals.
@@ 4.3 Ablation and Sensitivity Study
(a) Ablation Experiment Results
(b) Sensitivity Study Results
@@
We conduct ablation and parameter sensitivity studies to examine the contribution and robustness of each component in CRIB. The experiments are performed on PEMS-BAY dataset with four missing rates. In the Ablation Study (Fig. 4 (a)), we design three ablation experiments with configurations as follows: (1) w/o Uni-Atten: we replace the unified-variate attention mechanism with the vanilla attention mechanism. (2) w/o Consis: we remove the consistency regularization. (3) w/o IB: we remove the compactness and informativeness guidance of IB. In the Sensitivity Study (Fig. 4 (b)), we vary the weights assigned to the Embedding Size, IB weight: $\alpha$, and Consis Weight: $\gamma$ to study how each impacts model performance. We get observations as follows:
Obs. ❸: Capturing variate correlations and ensuring consistency are critical for direct forecasting. Both removing the unified-variate attention module (w/o Uni-Atten) and consistency regularization (w/o Consis) lead to a significant performance drop. This highlights the importance of modeling inter-variate dependencies to comprehend the true data correlations, especially when values are missing. Moreover, as shown in Tab. 2, consistency regularization is crucial for improving the model’s accuracy and stability, evidenced by lower prediction error and variance.
Obs. ❹: The Information Bottleneck principle is the model’s foundational component. The most severe performance degradation occurs when the IB guidance is removed (w/o IB). The relative stability of the full model and the other variants, contrasted with the sharp decline of the w/o IB variant, confirms that the IB principle is fundamental to our model’s ability to filter noise and achieve robust performance from incomplete data.
Obs. ❺: CRIB is robust to hyperparameter variations, though over-regularization can be detrimental under high missing rates. As shown in Fig. 4 (b), a larger embedding size generally correlates with better performance. However, the model remains effective even with a small embedding size (e.g., 32), demonstrating its efficiency in terms of computational and memory costs. For the IB and consistency regularization weights, we observe a trade-off. At low missing rates, higher weight values can improve accuracy. However, as the missing rate increases, excessively high weights tend to over-regularize the model, which can hinder its ability to capture complex variate correlations and thus degrade the final forecasting performance.
@@ 5 Related Work
The IB principle offers a framework for learning a compressed representation of an input that is maximally informative about a target task (Tishby et al., 2000). In time series, this is often implemented via Variational Autoencoders (VAEs) (Kingma, 2013; Voloshynovskiy et al., 2019). Existing methods like GP-VAE (Fortuin et al., 2020), MTS-IB (Ullmann et al., 2023), and RIB (Xu & Fekri, 2018) use the IB framework to model temporal dynamics. However, these approaches face a key limitation: a direct application of the IB principle can cause the model to concentrate too narrowly on observed features (Choi & Lee, 2023; Zhang et al., 2025b), thereby neglecting the broader variate correlations crucial for forecasting from incomplete data. In contrast to these works, our proposed CRIB applies the IB principle with a unified-attention mechanism and a consistency regularization, which encourages the model to capture stable representations and robust variate correlations even from sparse, incomplete inputs.
@@ Appendix A Full Derivation
The inequalities of the upper and lower bound in Eqs. 13 and 14 follow directly from the non-negativity of the KL-divergence and Entropy.
Appendix B Datasets
Table 3: Dataset Statistics.
@@ Appendix B Datasets
Statistics PEMS-BAY Metr-LA ETTh1 Electricity
Timesteps (T) 52116 34272 17420 26304
Variates (N) 325 207 7 321
Frequency 5 min 5 min 1 h 1 h
Mean Value 62.62 53.72 4.58 2538.79
Std Value 9.59 20.26 6.53 15027.57
We introduce information about datasets (Yu et al., 2024) as follows:
  • PEMS-BAY (Li et al., 2017): This is a traffic speed dataset collected by the California Transportation Agencies’ Performance Measurement System. It contains data collected by 325 sensors from January 1, 2017, to May 31, 2017. Each time series is sampled at a 5-minute interval, resulting in a total of 52,116 time slices.
  • METR-LA (Li et al., 2017): This is a traffic speed dataset collected using loop detectors located on the LA County road network. It contains data collected by 207 sensors from March 1, 2012, to June 30, 2012. Each time series is sampled at a 5-minute interval, resulting in a total of 34,272 time slices.
  • ETTh1 (Zhou et al., 2021): This is a dataset used for forecasting tasks, containing data from a power plant. It consists of measurements taken hourly, including features such as power consumption, temperature, and pressure. Each time series is sampled at a 1-hour interval, resulting in a total of 17,420 time slices.
  • Electricity (Wu et al., 2021): This dataset contains electricity consumption data. Each time series is sampled at a 1-hour interval, resulting in a total of 26,304 time slices.
Appendix C Baselines
  • BiTGraph (Chen et al., 2023): A model that jointly captures temporal correlations and spatial structures using biased Multi-Scale Instance PartialTCN and Biased GCN modules to effectively handle missing patterns in time series forecasting.
  • BRITS (Cao et al., 2018): A bidirectional RNN model that imputes missing values directly within a recurrent dynamical system, effectively handling correlations, nonlinear dynamics, and general missing data patterns.
  • GRIN (Cini et al., 2021): A graph neural network architecture designed for multivariate time series imputation, leveraging spatial and temporal message passing to reconstruct missing data.
  • SAITS (Du et al., 2023): A self-attention-based model for multivariate time series imputation that uses diagonally-masked self-attention blocks to capture temporal and feature correlations.
  • SPIN (Marisca et al., 2022): An attention-based spatial-temporal model for imputing multivariate time series, which avoids error propagation and does not rely on bidirectional encoding.
  • SegRNN (Lin et al., 2023): An RNN-based model using segment-wise iterations and parallel multi-step forecasting to reduce recurrence and improve accuracy, speed, and efficiency over Transformer baselines.
  • WPMixer (Murad et al., 2025): A MLP-based model (Wavelet Patch Mixer), leveraging the benefits of patching, multi-resolution wavelet decomposition, and mixing.
  • iTransformer (Liu et al., 2023): A restructured Transformer for time series forecasting that captures multivariate correlations via attention on variate tokens, enhancing performance and efficiency across variable lookback windows.
  • PatchTST (Nie et al., 2022): A Transformer-based model that segments time series into patches with a channel-independent design, enhancing long-term forecasting.
  • DLinear (Zeng et al., 2023): A model that uses a simple MLP as the predictor to forecast accurately and has achieved great success.
  • TimeXer (Wang et al., 2024b): A Transformer-based model that employs patch-level and variate-level representations respectively for endogenous and exogenous variables, with an endogenous global token as a bridge in-between.
  • PAttn (Tan et al., 2024): A simple Transformer-based model combining patching with one-layer attention.
Appendix D Full Experiments

### v3

@@ 3.5 Consistency Regularization
While the IB framework encourages learning a compact representation, high missing rates can still lead to unstable training as shown in App. E.2, where the model overfits to the specific variate in a given time window (Choi and Lee, 2023). To mitigate this and enhance robustness, we introduce a consistency regularization scheme (Bachman et al., 2014; Laine and Aila, 2016). The core intuition is that the model’s prediction should be invariant to the missingness. We achieve this by creating an augmented, more challenging view of the input, e.g, introducing additional noise to partially observed data. By enforcing that the representations learned from the observed and augmented views remain consistent, we regularize the model to handle missing values while stabilizing the refined representations instead of focusing excessively on a limited subset of observed data and neglecting crucial task-relevant variate correlations.
@@ 3.6 Model Learning
Table 2: Performance comparison on six additional datasets with a point missing pattern (average MAE and MSE across 20% to 70% missing rate). Best is Bold and second-best is Underlined.
Data Metric SegRNN WPMixer iTransformer PatchTST DLinear PAttn CSDI NeuralCDE ImputeFormer TimesNet Ours
ETTh2 MAE 0.163 0.171 0.225 0.182 0.225 0.211 0.519 0.320 0.152 0.182 0.141
MSE 0.061 0.066 0.124 0.090 0.117 0.106 0.684 0.188 0.054 0.076 0.048
ETTm1 MAE 0.339 0.381 0.398 0.377 0.505 0.389 0.658 0.498 0.400 0.361 0.313
MSE 0.526 0.637 0.671 0.606 0.967 0.638 1.284 0.853 0.601 0.565 0.453
ETTm2 MAE 0.130 0.153 0.206 0.182 0.217 0.190 0.527 0.234 0.123 0.170 0.116
MSE 0.041 0.054 0.113 0.092 0.107 0.098 0.681 0.110 0.037 0.065 0.034
Weather MAE 0.059 0.087 0.118 0.121 0.171 0.118 0.381 0.131 0.059 0.111 0.055
MSE 0.052 0.052 0.140 0.120 0.155 0.114 0.882 0.133 0.052 0.081 0.048
Exchange MAE 0.088 0.108 0.187 0.125 0.216 0.091 0.626 0.401 0.056 0.115 0.034
MSE 0.025 0.031 0.082 0.065 0.124 0.052 1.109 0.337 0.007 0.036 0.003
BeijingAir MAE 0.358 0.371 0.382 0.368 0.405 0.381 0.543 0.441 0.366 0.376 0.348
MSE 0.532 0.587 0.598 0.568 0.619 0.592 1.018 0.691 0.564 0.585 0.533
where $\alpha,\beta,\gamma\in\mathbb{R}$ are the preset balancing coefficients. This entire guidance helps CRIB extract the most important task-relevant information from the partially observed data while filtering out noise introduced by missing values.
4 Experiment
4.1 Experiment Settings
Datasets.
We evaluate our model on 11 MTSF datasets: PEMS-BAY, Metr-LA (Li et al., 2017), ETTh1, ETTh2, ETTm1, ETTm2 (Zhou et al., 2021), Weather (Miao et al., 2024), BeijingAir, Exchange (Lai et al., 2018), Electricity (Wu et al., 2021), and AQI (Yi et al., 2016). More information about these datasets is summarized in App. C.
Baselines.
We evaluate 16 representative baselines: (1) MTSF-M methods: BRITS (Cao et al., 2018), SAITS (Du et al., 2023), SPIN (Marisca et al., 2022), GRIN (Cini et al., 2021), and BiTGraph (Chen et al., 2023). (2) Imputation methods: CSDI (Tashiro et al., 2021), NeuralCDE (Kidger et al., 2020), ImputeFormer (Nie et al., 2024), and TimesNet (Wu et al., 2022). (3) Transformer-based MTSF methods: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), and PAttn (Tan et al., 2024). (4) MLP- and RNN-based MTSF methods: DLinear (Zeng et al., 2023), WPMixer (Murad et al., 2025), TimeXer (Wang et al., 2024b), and SegRNN (Lin et al., 2023).
Implementation Details.
Several forecasting baselines are not originally designed for MTSF-M settings. We therefore construct two-stage variants by combining them with SOTA imputation model TimesNet (Wu et al., 2022), which which imputes missing values before forecasting. To simulate a practical scenario where ground truth for missing values is unavailable during inference, TimesNet is trained with a 10% missing rate and then used to impute datasets with 20%, 40%, 60%, and 70% missing rates. The original models and the variants are denoted as Original and Imputed, respectively. More details are provided in App. D.1.
Training. We evaluate forecasting performance using Mean Squared Error (MSE) and Mean Absolute Error (MAE). Our training cost is in App. D.3.
4.2 Main Results
@@
The average performance comparisons are reported in  Tabs. 1 and 2, with full results and additional missing-pattern evaluations provided in App. H and Figure 3. We denote out-of-memory and improvement as OOM and IMP, respectively. Based on these results, we summarize our observations (Obs.) as follows:
Obs. ❶: CRIB demonstrates superior performance improvement in MTSF-M tasks. Overall, CRIB consistently improves forecasting accuracy over strong baselines across diverse datasets and missing-value settings. Specifically, CRIB reduces the MAE by over 18% on ETTh1 and over 13% on PEMS-BAY compared to the strongest baseline. We attribute this improvement to our model’s design, which integrates patch embedding, unified-variate attention, and consistency regularization under the IB principle, thus enabling CRIB to effectively filter noise from incomplete data while preserving essential predictive signals.
@@
4.3 Natural Missingness Evaluation
Table 3: Performance comparison on AQI datasets (Original vs. Imputed). The best results are highlighted in Bold, and the second-best is highlighted in Underline.
Data Metric SegRNN WPMixer iTransformer PatchTST DLinear PAttn CSDI NeuralCDE ImputeFormer TimesNet Ours
AQI_ORI MAE 0.604 0.624 0.608 0.627 0.598 0.621 0.858 0.798 0.795 0.648 0.555
MSE 0.804 0.843 0.818 0.844 0.741 0.843 1.448 1.438 1.313 0.925 0.663
AQI_IMP MAE 0.650 0.665 0.668 0.666 0.653 0.663 0.941 0.946 0.857 0.733 0.616
MSE 0.966 0.986 1.012 0.993 0.893 0.995 1.775 1.928 1.543 1.206 0.844
Obs. ❸: Imputation-based strategies consistently underperform direct prediction on naturally missing data. As shown in Tab. 3, models trained on the fully imputed dataset (AQI_IMP) yield higher error rates on observed entries compared to the original sparse data (AQI_ORI). We attribute this degradation to distributional perturbations introduced by unsupervised imputation. Lacking ground truth, the imputer generates synthetic artifacts that deviate from real dynamics. Consequently, the downstream model is forced to fit these erratic perturbations, which distorts feature representation and impairs its ability to capture authentic dependencies in valid observations. More implementation details are in App. D.2.
@@ 4.4 Ablation and Sensitivity Study
We further conduct ablation and parameter sensitivity studies on the PEMS-BAY, ETTh1, Electricity and Metr-LA datasets under four missing rates to examine the contribution and robustness of each component in CRIB. In the Ablation Study (Figure 4 (a) and App. F), we design three ablation experiments with configurations as follows: (1) w/o Uni-Atten: we replace the unified-variate attention mechanism with the vanilla attention mechanism. (2) w/o Consis: we remove the consistency regularization. (3) w/o IB: we remove the compactness and informativeness guidance of IB. We get observations as follows:
Obs. ❹: Capturing variate correlations and ensuring consistency are critical for direct forecasting. Removing either the unified-variate attention mechanism or the consistency regularization leads to clear performance degradation across all missing settings. This indicates that explicitly modeling inter-variate correlations and enforcing representation consistency are both essential for accurate forecasting under missingness. Moreover, as shown in Tab. 4, consistency regularization plays an important role in stabilizing training and reducing prediction variance, as reflected by the increased error and instability observed when it is removed.
Obs. ❺: The Information Bottleneck principle is the model’s foundational component. Among all ablation variants, eliminating the IB guidance results in the most severe performance drop. Compared with other components, the IB principle has a more fundamental impact on model behavior, as it directly governs the balance between filtering irrelevant noise and preserving predictive information from incomplete inputs. The sharp degradation of the w/o IB variant, contrasted with the relatively stable performance of other ablations, confirms that IB serves as the core mechanism that enables CRIB to learn robust and compact representations under missing values.
We additionally conduct the Sensitivity Study (Figure 4 (b) and App. G) of key hyperparameters, we vary the weights assigned to the Embedding Size, IB weight: $\alpha$, and Consis Weight: $\gamma$ to study how each impacts model performance.
Obs. ❻: CRIB is robust to hyperparameter variations, though over-regularization can be detrimental under high missing rates. As shown in Figure 4 (b), a larger embedding size generally correlates with better performance. However, the model remains effective even with a small embedding size (e.g., 32), demonstrating its efficiency in terms of computational and memory costs. For the IB and consistency regularization weights, we observe a trade-off. At low missing rates, higher weight values can improve accuracy. However, as the missing rate increases, excessively high weights tend to over-regularize the model, which can hinder its ability to capture complex variate correlations and thus degrade the final forecasting performance.
@@ 5 Related Work
The IB principle provides a theoretical framework for learning compact representations that retain maximal task-relevant information while discarding irrelevant variability (Tishby et al., 2000). In the context of time series modeling, IB is commonly implemented via Variational Autoencoders (VAEs) (Kingma, 2013; Voloshynovskiy et al., 2019). Existing methods like GP-VAE (Fortuin et al., 2020), MTS-IB (Ullmann et al., 2023), and RIB (Xu and Fekri, 2018) use the IB framework to model temporal dynamics. However, these approaches face a key limitation: a direct application of the IB principle can cause the model to concentrate too narrowly on observed features (Choi and Lee, 2023; Zhang et al., 2025b), thereby neglecting the broader variate correlations crucial for forecasting from incomplete data. In contrast to these works, our proposed CRIB applies the IB principle with a unified-attention mechanism and a consistency regularization. This design enables the model to leverage the regularization benefits of IB while avoiding excessive information collapse, resulting in representations that are both compact and robust to sparsity in the inputs.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Full Derivation
The inequalities of the upper and lower bound in Equations 13 and 14 follow directly from the non-negativity of the KL-divergence and Entropy.
Appendix B Quantitative Analysis of Introduction
Table 5: Performance comparison on PEMS-BAY dataset (40% Missing).
Metric CRIB BiTGraph DLinear TimesNet + DLinear TimeXer TimesNet + TimeXer
MAE 0.093 0.413 0.156 0.148 0.125 0.135
MSE 0.043 0.788 0.087 0.081 0.051 0.073
  • Detrimental Imputation: We demonstrate that imputation without ground truth is often harmful. As visualized in Figure 1, methods following the “imputation-then-prediction” paradigm fail to recover the true data distribution and instead reinforce biased patterns from partial observations.
  • Performance Degradation: We provide counter-intuitive evidence that imputation actively harms prediction accuracy. For instance, equipping the predictor TimeXer with the SOTA imputer TimesNet increases the MAE from 0.125 to 0.135 as shown in Tab. 5. It also has no help in understanding the data distribution and variate correlations as demonstrated in Figure 1.
Appendix C Datasets
To assess the model’s effectiveness and robustness in handling missing values, we introduce synthetic missingness by randomly removing data points at varying missing rates of 20%, 40%, 60%, and 70% with three different missing patterns. During the experiments, we normalized the data to facilitate better model fitting.
We introduce information about datasets in Tab. 6.
Table 6: Statistics of the 11 real-world datasets used in our experiments.
@@ Appendix C Datasets
Statistics ETTh1 ETTh2 ETTm1 ETTm2 Electricity PEMS-BAY Metr-LA BeijingAir Weather Exchange AQI
Time Steps 17,420 17,420 69,680 69,680 26,304 52,116 34,272 36,000 52,696 7,588 8759
Variates 7 7 7 7 321 325 207 7 21 8 36
Appendix D Implementation Details
D.1 Comparison Evaluations
We use Adam optimizer (Kingma, 2014) to learn the parameters of all models with $10^{-3}$ learning rate. The unified-variate attention of CRIB is configured with 2 layers and 4 heads, while the predictor is implemented as a simple 2-layer MLP. Both historical and future time window sizes are set to 24 for all methods, following the setting of BiTGraph (Chen et al., 2023). The patch length is set to 8, so every time series in a time window is patched into three tokens. The entire dataset is divided into training, validation, and testing sets with ratios of 60%, 20%, and 20%. All methods are trained on an Nvidia 4090 GPU with PyTorch 2.0. Hyperparameters of all baselines are consistent with their original papers.
D.2 Natural Missingness Evaluations
To fairly compare direct prediction with the imputation-then-prediction paradigm, we first pre-trained TimesNet (Wu et al., 2022) on the AQI training set for imputation by applying an additional 10% point-missing mask on observed entries, matching the dataset’s estimated natural missing rate. We then used the trained TimesNet to impute all NaNs in AQI and trained/evaluated downstream forecasting models on this fully imputed dataset (AQI_IMP). In contrast, we trained/evaluated the same models directly on the original AQI with natural missingness (AQI_ORI).
D.3 Training Cost
To ensure fair comparisons, we train all baseline models from scratch using identical dataset splits and experimental protocols, as detailed in Sec. 4.1. We evaluate computational efficiency by reporting the memory footprint and parameter counts of every model on ETTh1 as follows.
Conclusion:
The results demonstrate that CRIB maintains a computational cost comparable to efficient Transformer baselines (e.g., PatchTST (Nie et al., 2022)) while being more lightweight than complex methods such as CSDI (Tashiro et al., 2021) and TimesNet (Wu et al., 2022).
Table 7: Comparison of model efficiency in terms of parameter count and memory cost. CRIB achieves a balanced trade-off between performance and efficiency.
Model Parameters Memory (MB)
CRIB (Ours) 37,450 148.03
DLinear 1,200 18.99
PAttn 15,640 55.13
SegRNN 8,056 30.17
Transformer 57,063 189.38
iTransformer 39,768 154.18
PatchTST 41,528 179.80
TSMixer 6,837 21.16
WPMixer 44,370 50.16
CSDI 239,649 1,269.72
NeuralCDE 37,767 41.40
ImputeFormer 264,193 1,488.22
TimesNet 863,895 1,427.44
Appendix E Extra Experiments
E.1 Forecasting Results Visualization
We present a spatial visualization of forecasting results to demonstrate the effectiveness of CRIB under varying missing rates. Figure 5 shows the final timestamp in the historical time window and the first forecasting timestamp on the PEMS-BAY dataset. At lower missing rates (20% and 40%), by effectively leveraging inter-variate correlations extracted from the data, CRIB accurately predicts the future values. Even at higher missing rates (60% and 70%), CRIB can maintain stable performance and predict the spatial distribution of the PEMS-BAY datasets. These findings underscore CRIB’s capability to handle incomplete data and produce reliable predictions.
E.2 Unified-Variate Attention Maps Visualization
In Figure 6, we compare visualizations of directly applying IB on the Transformer with our proposed CRIB. In the first experiment, a transformer model serves as the predictor. The left two figures clearly show that directly applying IB to the model would force the model to focus on a few specific values (straight line attention), thereby neglecting global information. In contrast, the right figures reveal that CRIB can not only capture the original intra-variate temporal correlations in one attention head but also effectively uncovers cross-variate correlations in another, rather than relying solely on raw correlations. As a result, the final forecasting performance is improved remarkably by our unified-variate attention mechanism and consistency regularization scheme.
E.3 Experiments on Various Missing Patterns
Appendix F Extra Ablation Study
We have done an extra ablation study on three cases of loss of CRIB to prove its effectiveness. The ablation analysis across four datasets confirms the necessity of each component, as removing the Consistency Regularization ($\mathcal{L}_{\text{Consis}}$), Compactness ($\mathcal{L}_{\text{Comp}}$), or Informativeness ($\mathcal{L}_{\text{Pred}}$) objectives consistently leads to performance degradation, validating their collective role in ensuring robust and accurate forecasting in MTSF-M tasks.
Table 8: Ablation study on ETTh1 dataset.
ETTh1 0.2 0.4 0.6 0.7
MAE MSE MAE MSE MAE MSE MAE MSE
w/o Consis 0.237 0.206 0.274 0.270 0.338 0.413 0.402 0.584
w/o Reg 0.236 0.206 0.274 0.270 0.339 0.410 0.400 0.579
w/o Pred 0.432 0.559 0.505 0.698 0.655 1.066 0.796 1.486
Entire 0.220 0.171 0.251 0.249 0.267 0.296 0.288 0.361
Table 9: Ablation study on Elec dataset.
Elec 0.2 0.4 0.6 0.7
MAE MSE MAE MSE MAE MSE MAE MSE
w/o Consis 0.0201 0.0196 0.0267 0.0338 0.0346 0.0569 0.0415 0.0931
w/o Reg 0.0186 0.0175 0.0244 0.0286 0.0322 0.0544 0.0399 0.0980
w/o Pred 0.0881 0.5454 0.1243 0.9424 0.1818 1.8995 0.2283 2.9058
Entire 0.0150 0.0120 0.0230 0.0280 0.0300 0.0470 0.0380 0.0910
Table 10: Ablation study on Metr-LA dataset.
Metr-LA 0.2 0.4 0.6 0.7
MAE MSE MAE MSE MAE MSE MAE MSE
w/o Consis 0.271 0.332 0.257 0.275 0.271 0.306 0.314 0.366
w/o Reg 0.253 0.307 0.251 0.275 0.266 0.307 0.311 0.364
w/o Pred 0.442 0.418 0.624 0.529 0.985 1.177 1.276 1.959
Entire 0.248 0.271 0.249 0.272 0.265 0.305 0.309 0.356
Table 11: Ablation study on PEMS-BAY dataset.
PEMS-BAY 0.2 0.4 0.6 0.7
MAE MSE MAE MSE MAE MSE MAE MSE
w/o Consis 0.0959 0.0470 0.0882 0.0389 0.0976 0.0465 0.1296 0.0596
w/o Reg 0.0942 0.0453 0.0869 0.0373 0.0964 0.0456 0.1190 0.0591
w/o Pred 0.4006 0.2405 0.5618 0.3631 1.0832 1.2727 2.1480 4.8111
Entire 0.0830 0.0340 0.0850 0.0350 0.0930 0.0430 0.1100 0.0580
Appendix G Extra Sensitivity Study
We analyze hyperparameter sensitivity in Figure 4 (b) and conduct additional sensitivity studies. Empirical results indicate that the optimal settings are consistent across different datasets and missing rates. We set the weights for $\mathcal{L}_{\text{Comp}}$, $\mathcal{L}_{\text{Pred}}$, and $\mathcal{L}_{\text{Consis}}$ to $0^{-6}, 11$, and $1$ as default, respectively.
Table 12: Sensitivity analysis of $\mathcal{L}_{\text{Pred}}$ weight across all missing rates.
Weight 0% Missing 20% Missing 40% Missing 60% Missing 70% Missing
ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange
0.1 0.201 0.0186 0.243 0.0332 0.281 0.0439 0.348 0.0413 0.412 0.0373
0.5 0.198 0.0186 0.238 0.0301 0.273 0.0294 0.340 0.0362 0.406 0.0423
1.0 0.198 0.0186 0.237 0.0287 0.274 0.0253 0.337 0.0363 0.403 0.0517
2.0 0.199 0.0186 0.236 0.0261 0.273 0.0261 0.339 0.0363 0.400 0.0475
5.0 0.200 0.0186 0.237 0.0246 0.273 0.0258 0.338 0.0287 0.401 0.0320
Table 13: Sensitivity analysis of $\mathcal{L}_{\text{Comp}}$ weight across all missing rates.
Weight 0% Missing 20% Missing 40% Missing 60% Missing 70% Missing
ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange
10 0.214 0.0192 0.260 0.0509 0.314 0.0806 0.417 0.2131 0.506 0.2851
1 0.208 0.0188 0.251 0.0471 0.296 0.0835 0.365 0.1393 0.437 0.2293
$10^{-2}$ 0.199 0.0186 0.238 0.0300 0.274 0.0319 0.341 0.0435 0.406 0.0499
$10^{-6}$ 0.198 0.0186 0.237 0.0287 0.274 0.0253 0.337 0.0363 0.403 0.0517
$10^{-10}$ 0.199 0.0186 0.237 0.0276 0.273 0.0260 0.339 0.0417 0.403 0.0408
Table 14: Sensitivity analysis of $\mathcal{L}_{\text{Consis}}$ weight across all missing rates.
Weight 0% Missing 20% Missing 40% Missing 60% Missing 70% Missing
ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange ETTh1 Exchange
0.1 0.201 0.0186 0.237 0.0231 0.274 0.0270 0.340 0.0301 0.400 0.0341
0.5 0.199 0.0186 0.236 0.0260 0.274 0.0273 0.339 0.0378 0.401 0.0368
1.0 0.198 0.0186 0.237 0.0287 0.274 0.0253 0.337 0.0363 0.403 0.0517
2.0 0.199 0.0186 0.239 0.0306 0.275 0.0300 0.340 0.0352 0.404 0.0427
5.0 0.200 0.0186 0.240 0.0312 0.278 0.0377 0.344 0.0376 0.410 0.0357
Appendix H Full Experiments